OpenStudy (anonymous):

Which statement about the following equation is true? 2x2 – 9x + 2 = –1

OpenStudy (anonymous):

The discriminant is less than 0, so there are two real roots. The discriminant is less than 0, so there are two complex roots. The discriminant is greater than 0, so there are two real roots. The discriminant is greater than 0, so there are two complex roots.

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

will recieve fan and metal

OpenStudy (cwrw238):

2x2 – 9x + 2 = –1 2x^2 - 9x + 3 = 0 discriminant for ax^2 + bx + c = 0 is b^2-4ac so plug in a = 2, b = -9 and c = 3 into this formula and work it out

OpenStudy (cwrw238):

ok I'll start it off for you discriminant = (-9)^2 - 4*2*3

OpenStudy (anonymous):

alright ill get it ty

OpenStudy (anonymous):

i got 57

OpenStudy (anonymous):

so C

OpenStudy (cwrw238):

right so now you can find the correct option

OpenStudy (anonymous):

its greater than

OpenStudy (cwrw238):

yes

OpenStudy (cwrw238):

greater than 0

OpenStudy (anonymous):

So it would be C

OpenStudy (anonymous):

there are two real roots

OpenStudy (cwrw238):

right

OpenStudy (anonymous):

ty

OpenStudy (cwrw238):

no probs